OpenStudy (anonymous):

what is the procedure?? Log2(3-7x)=log2 (6x2)

OpenStudy (xapproachesinfinity):

hey! \[\log_{2} [3-7x]=\log_{2}[6x^2] \Longrightarrow 3-7x=6x^2 \]

OpenStudy (xapproachesinfinity):

we have the property: \[\log_{a} (b)=\log_{a}(c) \Longrightarrow b=c \]

OpenStudy (xapproachesinfinity):

now you just need to solve the quadratic equation \(\large 6x^2+7x-3=0\)

OpenStudy (anonymous):

ok thanks

OpenStudy (xapproachesinfinity):

wlc!

OpenStudy (xapproachesinfinity):

just try the quadratic formula